### v1

# Contrast: A Hybrid Architecture of Transformers and State Space Models for Low-Level Vision

Abstract
Transformers have become increasingly popular for image super-resolution (SR) tasks due to their strong global context modeling capabilities. However, their quadratic computational complexity necessitates the use of window-based attention mechanisms, which restricts the receptive field and limits effective context expansion. Recently, the Mamba architecture has emerged as a promising alternative with linear computational complexity, allowing it to avoid window mechanisms and maintain a large receptive field. Nevertheless, Mamba faces challenges in handling long-context dependencies when high pixel-level precision is required, as in SR tasks. This is due to its hidden state mechanism, which can compress and store a substantial amount of context but only in an approximate manner, leading to inaccuracies that transformers do not suffer from. In this paper, we propose Contrast, a hybrid SR model that combines Convolutional, Transformer, and State Space components, effectively blending the strengths of transformers and Mamba to address their individual limitations. By integrating transformer and state space mechanisms, Contrast compensates for the shortcomings of each approach, enhancing both global context modeling and pixel-level accuracy. We demonstrate that combining these two architectures allows us to mitigate the problems inherent in each, resulting in improved performance on image super-resolution tasks.
1 Introduction
Figure 1: Model comparison on the Urban100 dataset for $\times$4 SR. The plot illustrates the trade-off between PSNR and parameter count, where higher PSNR values and fewer parameters (upper-left) indicate better performance. This highlights the effectiveness of Contrast on the Urban100 benchmark.
Single image super-resolution (SR) aims to reconstruct a high-resolution (HR) image from a low-resolution (LR) input, serving as a fundamental task in low-level vision. This problem is inherently ill-posed, as multiple plausible solutions can exist for any given LR input. In recent years, a variety of methods have been explored to improve SR performance, balancing computational efficiency with the ability to capture fine-grained details.
Convolutional Neural Networks (CNNs) [15, 16, 33] were the early choice for SR tasks [10, 8, 27, 45], effectively capturing local structures in images. However, due to the local nature of convolutions, CNNs struggle to model long-range dependencies, a key component for capturing contextual information in complex images. This limitation has prompted the search for alternative methods that can better handle global context in SR.
Transformers [38], originally developed for natural language processing (NLP), have shown strong potential in computer vision, particularly for high-level tasks [11, 36, 37], due to their self-attention (SA) mechanism, which directly models long-range dependencies. By capturing global context, transformers help overcome the local limitations of CNNs. However, their quadratic computational complexity with image resolution becomes prohibitive in high-resolution SR tasks. To mitigate this, SR models typically employ window-based self-attention [24, 19], limiting the attention mechanism to local windows to reduce computational costs. While this approach effectively addresses the quadratic complexity, it introduces a constrained receptive field that is difficult to expand, limiting the model’s ability to capture broader image context essential for high-quality SR outcomes.
Recently, the Mamba architecture [12, 9] has emerged as a promising alternative, emphasizing efficient global modeling with linear complexity scaling. Mamba is capable of extending receptive fields across the entire image without relying on window mechanisms, thereby alleviating issues associated with limited receptive fields in window-based attention. This makes it particularly attractive for SR tasks, where capturing global context is crucial. Furthermore, increasing Mamba’s efficiency does not require the introduction of window mechanisms, allowing it to maintain a global receptive field inherently.
However, Mamba faces challenges when high pixel-level precision is required, as in SR tasks. Its hidden state mechanism allows it to compress and store substantial context information, but the representation is approximate rather than exact. This approximation can lead to inaccuracies in modeling long-context dependencies, especially when fine details are essential for accurate image reconstruction. In contrast, transformers do not suffer from this problem due to their precise attention mechanisms.
In this work, we demonstrate that by combining the strengths of transformers and Mamba, we can address the limitations inherent in each architecture. We propose Contrast, a hybrid SR model that integrates Convolutional, Transformer, and State Space components, effectively blending the advantages of both transformers and Mamba. The hybrid Contrast model leverages the receptive field expansion and efficient complexity scaling of Mamba, while simultaneously benefiting from the transformer’s capacity to model complex spatial dependencies with high pixel-level accuracy. By merging these capabilities, Contrast enhances the modeling of both local and global context, providing a balanced solution for SR that neither architecture achieves alone.
Figure 1 demonstrates that Contrast outperforms other models on the Urban100 [17] dataset in terms of PSNR relative to the number of parameters. Urban100 comprises high-resolution images with repetitive patterns, posing a challenge ideal for models with extensive receptive fields. Notably, Contrast achieves a PSNR of 27.92 with 14.1 million parameters, compared to MambaIR (ECCV’24) [14] which attains 27.68 PSNR with 16.7 million parameters, and HAT (CVPR’23) [3] which reaches 27.97 PSNR with 20.8 million parameters. This highlights the efficiency of Contrast, a hybrid model combining transformer and Mamba architectures in a ratio of one transformer block to six Mamba blocks, in delivering high-quality image reconstruction with fewer parameters, especially when contrasted with MambaIR and HAT.
In summary, adding window mechanisms to transformers can significantly alleviate the issue of quadratic complexity but at the cost of a limited receptive field that is challenging to expand. Mamba, with its linear complexity, avoids the need for window mechanisms altogether, inherently possessing a global receptive field. By integrating the two architectures, Contrast effectively compensates for their respective shortcomings, achieving superior performance in image super-resolution tasks.
Figure 2: The network architecture of our method. (a) The architecture of Contrast. (b) Illustration of VSS Block. (c) Illustration of Overlapping Cross-Attention Block (OCAB).
2 Related Work
Image Super-Resolution. Deep learning has revolutionized image super-resolution (SR), with Convolutional Neural Networks (CNNs) initially setting the benchmark for performance. Pioneering work like SRCNN [10] introduced CNNs to SR, achieving significant improvements over traditional methods. Advanced architectures such as RCAN [45] utilized deep residual networks exceeding 400 layers to enhance feature extraction. Additionally, attention mechanisms [45, 40] were integrated to focus on important spatial and channel-wise information. Despite these advancements, CNNs inherently struggle to capture long-range dependencies due to the local nature of convolution operations, limiting their ability to model global context effectively.
Vision Transformers in Low-Level Vision. Transformers [38], originally designed for natural language processing, have shown great promise in computer vision tasks [11, 36, 37] due to their self-attention mechanism, which can model long-range dependencies. In SR, Transformers help overcome the limitations of CNNs by capturing global context. However, the quadratic computational complexity of self-attention with respect to input size poses significant challenges for high-resolution images typical in SR tasks. To alleviate this, models like SwinIR [24, 19] employ window-based self-attention, where attention is calculated within local windows to reduce computational costs. While this approach effectively mitigates quadratic complexity, it introduces a limited receptive field that is difficult to expand, constraining the model’s ability to capture broader image context essential for high-quality SR outcomes.
State Space Models. State space models (SSMs) have recently emerged as a promising alternative for modeling long-range dependencies with linear computational complexity. Architectures like Mamba [12, 9] exploit SSMs to efficiently capture global context, inherently providing a global receptive field. This makes them particularly attractive for SR tasks, where understanding the entire image context is crucial. However, SSM-based models like Mamba face challenges when precise pixel-level accuracy is required. The hidden state mechanism compresses context information in an approximate manner, which can lead to inaccuracies in modeling long-context dependencies. Additionally, their sequential scanning mechanism makes it difficult to capture diagonal dependencies, which are important for reconstructing detailed spatial patterns.
Hybrid Models. Numerous architectures have explored hybrid models that integrate Transformers with SSMs like Mamba, often employing full attention mechanisms in the Transformer components [39]. However, few studies have delved into providing a comprehensive explanation for the effectiveness of such integrations. Our work distinguishes itself by offering greater clarity and laying the groundwork for the development of these hybrid ideas. Specifically, our proposed Contrast model combines Transformer blocks with windowed attention mechanisms and Mamba blocks to harness their respective strengths. The Mamba blocks provide a global receptive field, effectively capturing long-range dependencies essential for super-resolution tasks. In parallel, Transformer blocks with windowed attention refine the global information from the Mamba blocks, enhancing the model’s ability to capture detailed features. Additionally, we integrate convolutional layers within the MLP layers to model local pixel relationships, particularly focusing on diagonal pixels. This convolutional enhancement ensures that Contrast maintains high pixel-level accuracy while effectively capturing both local and global dependencies. By synergizing the global context capabilities of Mamba with the refinement capabilities of Transformer blocks and enhancing local pixel relationships through convolutional layers, Contrast overcomes the individual limitations of Transformers and SSMs when used in isolation.
3 Methodology
3.1 Architecture
The Contrast model is designed as a hybrid framework with three primary modules: shallow feature extraction, deep feature extraction, and image reconstruction, as illustrated in Fig. 2. This architecture balances computational efficiency and performance in super-resolution (SR) tasks by combining Visual State Space (VSS) Blocks[23], Overlapping Cross-Attention Blocks (OCAB)[3], and Spatial Gated Feed-Forward Networks (SGFN)[5]. Each module is tailored to extract and refine features at multiple levels, progressively enhancing the image from low resolution to high resolution.
3.1.1 Shallow Feature Extraction
Given a low-resolution (LR) input image $I_{LR}\in\mathbb{R}^{H\times W\times 3}$, the shallow feature extraction module first applies a convolutional layer to map $I_{LR}$ to an initial feature space:
$F_{S}=\text{Conv}(I_{LR}),$ (1)
where $F_{S}\in\mathbb{R}^{H\times W\times C}$ represents the shallow feature map, with $H$ and $W$ denoting the spatial dimensions of the input image, and $C$ the number of feature channels. This initial step provides a feature-rich representation for further processing.
3.1.2 Deep Feature Extraction
The deep feature extraction module builds upon the shallow features $F_{S}$ to capture complex spatial and channel relationships, outputting a deeper feature representation $F_{D}\in\mathbb{R}^{H\times W\times C}$. This module is composed of $N_{1}$ stacked Residual Groups (RGs), each containing $N_{2}$ Visual State Space (VSS) Blocks followed by an Overlapping Cross-Attention Block (OCAB).
In each Residual Group (RG), features are processed through a sequence of VSS Blocks, arranged to optimize computational efficiency and performance in hybrid architectures. Specifically, we utilize a structure of six Mamba blocks followed by a single Transformer block, a ratio that has demonstrated effectiveness and was empirically validated in prior work [39]. This setup aligns with the original HAT architecture, where we replaced the (S)W-MSA modules with Mamba blocks while preserving the original six-to-one block ratio. Although we considered adding more Transformer layers, the Overlapping Cross-Attention Block (OCAB) required for cross-window information flow is computationally intensive due to its overlapping nature. This increased the training time substantially and slowed the model to a point where further improvements in metrics could not justify the trade-off in efficiency, prompting us to maintain the initial six-to-one configuration.
The nested structure of the deep feature extraction module, composed of $N_{1}$ RGs, progressively refines the shallow features $F_{S}$. The output of each RG can be represented as:
$F_{RG}^{(j)}=\text{Conv}(\text{OCAB}(\text{VSS}_{N_{2}}(F_{RG}^{(j-1)})))+F_{RG}^{(j-1)},$ (2)
where $j$ denotes the index of the current RG, $F_{RG}^{(j-1)}$ is the input to the RG, and $\text{VSS}_{N_{2}}$ represents the sequential application of $N_{2}$ VSS Blocks within the RG. Each VSS Block applies a state-space model to capture global dependencies with reduced computational overhead, as follows:
$F_{VSS}^{(i)}=\text{VSS}(F_{VSS}^{(i-1)}),$ (3)
where $i$ denotes the index of the VSS Block within an RG.
Following the VSS Blocks, the OCAB operation integrates cross-window information to enhance spatial coherence, ensuring a more continuous representation across feature windows. This is formulated as:
$F_{OCAB}=\text{OCAB}(F_{VSS}^{(N_{2})}),$ (4)
where $F_{OCAB}$ represents the output after the OCAB step, which is then refined by a convolution layer to produce the final output of the RG.
Thus, after passing through all $N_{1}$ Residual Groups, the deep feature extraction module yields the deep feature representation $F_{D}$, where each RG applies the combined operations of VSS Blocks, OCAB, and residual connections to progressively enrich the feature map.
3.1.3 Spatial Gated Feed-Forward Network (SGFN)
[…]
$\begin{gathered}\hat{X^{\prime}}=\sigma(W_{p}^{1}\hat{X}),\quad[\hat{X_{1}^{\prime}},\hat{X_{2}^{\prime}}]=\hat{X^{\prime}},\\ {\rm SGFN}(\hat{X})=W_{p}^{2}(\hat{X_{1}^{\prime}}\odot(W_{d}\hat{X_{2}^{\prime}})),\end{gathered}$ (5)
where $W_{p}^{1}$ and $W_{p}^{2}$ indicate linear projection layers, $\sigma$ is the GELU activation function, and $W_{d}$ represents the learnable parameters for the depth-wise convolution. Both $\hat{X_{1}^{\prime}}$ and $\hat{X_{2}^{\prime}}$ lie in $\mathbb{R}^{H\times W\times\frac{C^{\prime}}{2}}$ space, with $C^{\prime}$ denoting SGFN’s hidden dimension. Unlike the standard FFN, SGFN is designed to capture non-linear spatial information and reduce channel redundancy in fully-connected layers.
3.1.4 Image Reconstruction
The image reconstruction module takes the deep feature representation $F_{D}$ and upscales it to produce the final high-resolution (HR) output image $I_{HR}\in\mathbb{R}^{H_{out}\times W_{out}\times 3}$. The upsampling is performed using the pixel shuffle method [32], ensuring efficient spatial upscaling:
$F_{up}=\text{PixelShuffle}(F_{D}),$ (6)
where $F_{up}$ is the upscaled feature map. After upsampling, a convolutional layer aggregates the upscaled features to produce the HR image:
$I_{HR}=\text{Conv}(F_{up}).$ (7)
This series of transformations enhances the input LR image into a high-quality SR output, maintaining spatial continuity and ensuring fine-grained detail preservation across the reconstructed image.
4 Experiments
4.1 Experimental Settings
Implementation Details. We build two variants of the Contrast model with different complexity levels, called Contrast and Contrast-S. For the main Contrast model, we use 6 Residual Groups (RGs) with an embedding dimension of 210, a window size of 32, an MLP ratio of 2, an SSM state dimension of 1, and an SSM ratio of 1, as shown in the VMamba architecture. In the Contrast-S variant, we also use 6 RGs but with a reduced embedding dimension of 150 and a smaller window size of 16, while keeping the MLP ratio, SSM state dimension, and SSM ratio the same as in the main model.
In both models, we use the Mamba-1 module rather than Mamba-2. The authors of Mamba [12, 9] have noted that Mamba-2 provides minimal improvements in metrics over Mamba-1, with the primary enhancements focused on speeding up large language models (LLMs). For our super-resolution (SR) task, where we employ smaller embedding dimensions, Mamba-2 performed significantly slower than Mamba-1 in our experiments. Thus, Mamba-1 was chosen for Contrast and Contrast-S to balance efficiency and performance in SR.
Data and Evaluation. We follow the standard practices established in previous works [4, 5] for training and evaluating our models. Specifically, we use two large-scale datasets for training: DIV2K [35] and Flickr2K [21]. For evaluation, we test on five widely-used benchmark datasets: Set5 [2], Set14 [42], B100 [25], Urban100 [17], and Manga109 [26].
Our experiments cover upscaling factors of $\times$2, $\times$3, and $\times$4. Low-resolution (LR) images are generated from high-resolution (HR) images using bicubic degradation. To evaluate super-resolution (SR) performance, we use Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [41], both calculated on the Y channel (luminance) of the YCbCr color space. These metrics provide objective assessments of image quality and structural preservation in SR results.
Training Settings. We train the models with a patch size of 64×64 and a batch size of 32 for 500K iterations. The optimization is performed by minimizing the L1 loss using the Adam optimizer [18] with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. The initial learning rate is set to $1\times 10^{-4}$ for the main Contrast model and $2\times 10^{-4}$ for Contrast-S. The learning rate is halved at milestones: [250K, 400K, 450K, 475K]. During training, data augmentation is applied by randomly rotating images by 90°, 180°, and 270°, along with horizontal flips. Our model is implemented in PyTorch[31] and Triton[34] and trained on 4 A6000 GPUs.
4.2 Ablation Study
In this section, we conduct a series of ablation studies to investigate the effectiveness of different components in our proposed Contrast model. We aim to understand how the integration of Transformer and State Space components contributes to the model’s performance, and we analyze the impact of architectural choices on training dynamics and final results. All experiments are conducted with models of comparable parameter sizes to ensure a fair comparison.
Effect of Replacing Transformer Blocks with SS2D. We begin by exploring the impact of replacing the (Shifted) Window Multi-Head Self-Attention ((S)W-MSA) blocks in the baseline Transformer model HAT [3] with the State Space 2D (SS2D) blocks from Mamba [12]. This modification results in an initial hybrid architecture, allowing us to evaluate whether SS2D provides a more effective solution for capturing dependencies and spatial context compared to traditional self-attention mechanisms.
Training Dynamics of Pure Mamba, Transformer, and Hybrid Models. To further investigate the individual contributions of the Transformer and Mamba architectures, we trained three models:
  • Pure Mamba: A model where all Transformer blocks are replaced with SS2D blocks, and Channel Attention Blocks (CABs) are removed.
  • Pure Transformer: The baseline HAT model without modifications.
  • Hybrid Contrast: The baseline HAT model with (S)W-MSA replaced by SS2D blocks.
All models have approximately the same number of parameters to ensure a fair comparison. We observed an interesting dynamic during training, as illustrated in Figure 3. The pure Mamba model achieves high PSNR values early in training but exhibits slow progress afterward. In contrast, the Transformer model starts with lower PSNR values but continues to improve steadily, eventually surpassing the pure Mamba model. The hybrid Contrast model combines the strengths of both architectures: it starts with high PSNR values like Mamba and continues to improve steadily like the Transformer, ultimately outperforming both.
Figure 3: PSNR (dB) during training of pure Mamba [12], pure Transformer (HAT [3]), and our proposed Hybrid Contrast model. Results are evaluated on the Urban100 [17] dataset for $\times$4 SR.
This training behavior is consistently observed across all five datasets: Set5, Set14, B100, Urban100, and Manga109. The hybrid model’s ability to combine the rapid initial learning of Mamba with the sustained improvement of the Transformer suggests that the integration of SS2D blocks with Transformer components effectively leverages the advantages of both architectures.
Impact of Channel Attention Blocks. We also investigated the role of Channel Attention Blocks (CABs) in our hybrid model. CABs were originally included in the baseline model, but we wanted to assess whether they are essential in our hybrid architecture or if removing them could lead to a more efficient model without compromising performance. To this end, we compared two versions of the hybrid model:
  • With CAB: The hybrid model including CABs.
  • Without CAB: The hybrid model without CABs, with additional layers and adjusted embedding dimensions to compensate for the removal and maintain an equal number of parameters.
The results, presented in Table 1, indicate that removing the CABs not only simplifies the model but also leads to improved PSNR values across all evaluated datasets. This suggests that the hybrid model benefits more from increased depth and embedding capacity than from the inclusion of CABs.
Table 1: Ablation study on the impact of Channel Attention Blocks (CABs). We compare models with and without CABs, adjusting the number of layers and embedding dimensions to keep the parameter count approximately equal. Results are reported in PSNR (dB) for $\times$4 SR.
Model Set5 Set14 B100 Urban100 Manga109
With CAB 32.44 28.78 27.68 26.57 31.06
Without CAB 32.53 28.88 27.69 26.65 31.16
Replacing MLP with SGFN. We further explored the effect of replacing the standard Multi-Layer Perceptron (MLP) layers in our model with the Spatial Gated Feed-forward Network (SGFN). The SGFN is designed to enhance the model’s ability to capture spatial information by integrating gating mechanisms. We trained two versions of the model:
  • With MLP: The hybrid model using standard MLP layers.
  • With SGFN: The hybrid model with MLP layers replaced by SGFN.
The results, shown in Table 2, demonstrate that using SGFN leads to improved performance on most datasets.
Table 2: Comparison of models using MLP and SGFN layers. Results are reported in PSNR (dB) for $\times$4 SR.
Model Set5 Set14 B100 Urban100 Manga109
With MLP 32.44 28.84 27.73 26.72 31.22
With SGFN 32.50 28.87 27.73 26.74 31.24
LAM Visualization Analysis. To better understand how each model captures dependencies and spatial context, we utilize local attribution map (LAMs) [13]. Figure 4 presents the LAM visualizations for pure Mamba, pure Transformer, and our hybrid Contrast model.
Figure 4: LAM visualizations and prediction results. The first and second rows show LAM visualizations for MambaIR [14], the Transformer model, and our hybrid Contrast model. The third row presents the corresponding SR predictions. The red square on the HR image marks the region used for LAM calculation. MambaIR exhibits a more global receptive field with a preference for features aligned horizontally and vertically, while the Transformer focuses more locally. Our Contrast model effectively combines these characteristics, capturing both global context and precise spatial details.
From the LAMs, we observe that MambaIR has a global receptive field, but with a tendency to emphasize horizontal and vertical features over diagonal ones. The Transformer model shows a more localized attention pattern. Our hybrid Contrast model successfully captures a broad receptive field while maintaining focus on important spatial regions, especially those with higher resolution or closer to the camera. This visualization underscores the advantage of our hybrid approach in balancing global context modeling and precise spatial attention, supporting our quantitative findings.
Set5 Set14 B100 Urban100 Manga109
Method Scale PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
EDSR [20] $\times$2 38.11 0.9602 33.92 0.9195 32.32 0.9013 32.93 0.9351 39.10 0.9773
RCAN [45] $\times$2 38.27 0.9614 34.12 0.9216 32.41 0.9027 33.34 0.9384 39.44 0.9786
SAN [7] $\times$2 38.31 0.9620 34.07 0.9213 32.42 0.9028 33.10 0.9370 39.32 0.9792
RFANet [22] $\times$2 38.26 0.9615 34.16 0.9220 32.41 0.9026 33.33 0.9389 39.44 0.9783
HAN [30] $\times$2 38.27 0.9614 34.16 0.9217 32.41 0.9027 33.35 0.9385 39.46 0.9785
CSNLN [28] $\times$2 38.28 0.9616 34.12 0.9223 32.40 0.9024 33.25 0.9386 39.37 0.9785
NLSA [29] $\times$2 38.34 0.9618 34.08 0.9231 32.43 0.9027 33.42 0.9394 39.59 0.9789
ELAN [44] $\times$2 38.36 0.9620 34.20 0.9228 32.45 0.9030 33.44 0.9391 39.62 0.9793
DFSA [1] $\times$2 38.38 0.9620 34.33 0.9232 32.50 0.9036 33.66 0.9412 39.98 0.9798
SwinIR [19] $\times$2 38.42 0.9623 34.46 0.9250 32.53 0.9041 33.81 0.9427 39.92 0.9797
Swin2SR [6] $\times$2 38.43 0.9623 34.48 0.9256 32.54 0.9050 33.89 0.9431 39.88 0.9798
ART [43] $\times$2 38.56 0.9629 34.59 0.9267 32.58 0.9048 34.30 0.9452 40.24 0.9808
CAT-A [4] $\times$2 38.51 0.9626 34.78 0.9265 32.59 0.9047 34.26 0.9440 40.10 0.9805
DAT [5] $\times$2 38.58 0.9629 34.81 0.9272 32.61 0.9051 34.37 0.9458 40.33 0.9807
SRFormer [46] $\times$2 38.51 0.9627 34.44 0.9253 32.57 0.9046 34.09 0.9449 40.07 0.9802
HAT [3] $\times$2 38.63 0.9630 34.86 0.9274 32.62 0.9053 34.45 0.9466 40.26 0.9809
MambaIR [14] $\times$2 38.57 0.9627 34.67 0.9261 32.58 0.9048 34.15 0.9446 40.28 0.9806
Contrast-S (ours) $\times$2 38.49 0.9624 34.58 0.9256 32.54 0.9041 34.06 0.9438 40.06 0.9802
Contrast (ours) $\times$2 38.58 0.9630 34.68 0.9261 32.57 0.9046 34.45 0.9467 40.30 0.9808
EDSR [20] $\times$3 34.65 0.9280 30.52 0.8462 29.25 0.8093 28.80 0.8653 34.17 0.9476
RCAN [45] $\times$3 34.74 0.9299 30.65 0.8482 29.32 0.8111 29.09 0.8702 34.44 0.9499
SAN [7] $\times$3 34.75 0.9300 30.59 0.8476 29.33 0.8112 28.93 0.8671 34.30 0.9494
RFANet [22] $\times$3 34.79 0.9300 30.67 0.8487 29.34 0.8115 29.15 0.8720 34.59 0.9506
HAN [30] $\times$3 34.75 0.9299 30.67 0.8483 29.32 0.8110 29.10 0.8705 34.48 0.9500
CSNLN [28] $\times$3 34.74 0.9300 30.66 0.8482 29.33 0.8105 29.13 0.8712 34.45 0.9502
NLSA [29] $\times$3 34.85 0.9306 30.70 0.8485 29.34 0.8117 29.25 0.8726 34.57 0.9508
ELAN [44] $\times$3 34.90 0.9313 30.80 0.8504 29.38 0.8124 29.32 0.8745 34.73 0.9517
DFSA [1] $\times$3 34.92 0.9312 30.83 0.8507 29.42 0.8128 29.44 0.8761 35.07 0.9525
SwinIR [19] $\times$3 34.97 0.9318 30.93 0.8534 29.46 0.8145 29.75 0.8826 35.12 0.9537
ART [43] $\times$3 35.07 0.9325 31.02 0.8541 29.51 0.8159 30.10 0.8871 35.39 0.9548
CAT-A [4] $\times$3 35.06 0.9326 31.04 0.8538 29.52 0.8160 30.12 0.8862 35.38 0.9546
DAT [5] $\times$3 35.16 0.9331 31.11 0.8550 29.55 0.8169 30.18 0.8886 35.59 0.9554
SRFormer [46] $\times$3 35.02 0.9323 30.94 0.8540 29.48 0.8156 30.04 0.8865 35.26 0.9543
HAT [3] $\times$3 35.07 0.9329 31.08 0.8555 29.54 0.8167 30.23 0.8896 35.53 0.9552
MambaIR [14] $\times$3 35.08 0.9323 30.99 0.8536 29.51 0.8157 29.93 0.8841 35.43 0.9546
Contrast-S (ours) $\times$3 35.02 0.9322 30.97 0.8541 29.49 0.8151 29.93 0.8838 35.25 0.9538
Contrast (ours) $\times$3 35.06 0.9324 31.00 0.8541 29.51 0.8158 30.17 0.8884 35.45 0.9549
EDSR [20] $\times$4 32.46 0.8968 28.80 0.7876 27.71 0.7420 26.64 0.8033 31.02 0.9148
RCAN [45] $\times$4 32.63 0.9002 28.87 0.7889 27.77 0.7436 26.82 0.8087 31.22 0.9173
SAN [7] $\times$4 32.64 0.9003 28.92 0.7888 27.78 0.7436 26.79 0.8068 31.18 0.9169
RFANet [22] $\times$4 32.66 0.9004 28.88 0.7894 27.79 0.7442 26.92 0.8112 31.41 0.918
HAN [30] $\times$4 32.64 0.9002 28.90 0.7890 27.80 0.7442 26.85 0.8094 31.42 0.9177
CSNLN [28] $\times$4 32.68 0.9004 28.95 0.7888 27.80 0.7439 27.22 0.8168 31.43 0.9201
NLSA [29] $\times$4 32.59 0.9000 28.87 0.7891 27.78 0.7444 26.96 0.8109 31.27 0.9184
ELAN [44] $\times$4 32.75 0.9022 28.96 0.7914 27.83 0.7459 27.13 0.8167 31.68 0.9226
DFSA [1] $\times$4 32.79 0.9019 29.06 0.7922 27.87 0.7458 27.17 0.8163 31.88 0.9266
SwinIR [19] $\times$4 32.92 0.9044 29.09 0.7950 27.92 0.7489 27.45 0.8254 32.03 0.9260
Swin2SR [6] $\times$4 32.92 0.9039 29.06 0.7946 27.92 0.7505 27.51 0.8271 31.03 0.9256
ART [43] $\times$4 33.04 0.9051 29.16 0.7958 27.97 0.7510 27.77 0.8321 32.31 0.9283
CAT-A [4] $\times$4 33.08 0.9052 29.18 0.7960 27.99 0.7510 27.89 0.8339 32.39 0.9285
DAT [5] $\times$4 33.08 0.9055 29.23 0.7973 28.00 0.7515 27.87 0.8343 32.51 0.9291
SRFormer [46] $\times$4 32.93 0.9041 29.08 0.7953 27.94 0.7502 27.68 0.8311 32.21 0.9271
HAT [3] $\times$4 33.04 0.9056 29.23 0.7973 28.00 0.7517 27.97 0.8368 32.48 0.9292
MambaIR [14] $\times$4 33.03 0.9046 29.20 0.7961 27.98 0.7503 27.68 0.8287 32.32 0.9272
Contrast-S (ours) $\times$4 32.92 0.9038 29.11 0.7947 27.93 0.7489 27.59 0.8280 32.21 0.9264
Contrast (ours) $\times$4 32.94 0.9030 29.20 0.7973 27.98 0.7508 27.92 0.8357 32.38 0.9283
Contrast+ (ours) $\times$4 33.09 0.9055 29.23 0.7980 28.00 0.7520 28.07 0.8385 32.61 0.9299
Table 3: Quantitative comparison with state-of-the-art methods. The best and second-best results are coloured red and blue.
4.3 Comparison with State-of-the-Art Methods
We compare our proposed models, Contrast-S and Contrast, with several state-of-the-art (SOTA) image super-resolution methods. Consistent with prior studies [45, 19], we employ a self-ensemble strategy during testing, denoted by the symbol “$+$”. The quantitative results are presented in Table 3, and visual comparisons are provided in Figure 5.
Figure 5: Visual comparison for $\times$4 SR on the Urban100 dataset. Our Contrast model produces sharper and more detailed images, effectively reconstructing fine textures and patterns.
Quantitative Results. As shown in Table 3, our models achieve competitive performance, particularly excelling on the Urban100 [17] dataset. Our Contrast model attains a PSNR of 27.92 dB on Urban100, closely approaching the performance of HAT [3], which achieves 27.97 dB but with approximately 50% more parameters. Compared to MambaIR [14], which is based on the Mamba architecture we aimed to enhance, our Contrast model surpasses it by 0.24 dB on Urban100 while using fewer parameters. Our smaller model, Contrast-S, also performs comparably, trailing MambaIR by only 0.09 dB despite having nearly half the number of parameters.
It’s important to note that our primary focus was on enhancing performance for datasets like Urban100, which contain high-resolution images with complex structures and repetitive patterns. These characteristics make Urban100 particularly challenging and ideal for models capable of capturing extensive global context. Our hybrid approach effectively addresses this, leveraging Mamba’s global receptive field and the Transformer’s precise contextual modeling to achieve superior results on Urban100.
On other datasets such as Set5, Set14, and B100, which consist of low-resolution images without significant patterns or contain low-quality photographs, our models do not outperform all existing methods. This outcome aligns with our expectations, as these datasets often favor deeper networks with larger parameter counts that can capture fine-grained details through extensive depth. Our goal was not to surpass SOTA results across all benchmarks but to demonstrate the effectiveness of our hybrid model in scenarios where capturing global context is crucial.
Visual Results. Figure 5 showcases visual comparisons on challenging samples from the Urban100 dataset. Our Contrast model effectively reconstructs intricate details and repetitive patterns, such as building facades and window grids, providing sharper and more detailed images than both MambaIR and other Transformer-based models.
5 Conclusion
In this paper, we proposed Contrast, a novel hybrid super-resolution model that effectively combines convolutional layers, Transformer blocks, and state space components. By integrating the global receptive field of Mamba with the precise contextual modeling capabilities of Transformers, our model addresses the limitations inherent in each individual approach.
Our competitive performance on the Urban100 dataset validates this hybrid approach, demonstrating that Contrast can effectively capture complex spatial patterns and deliver high-quality reconstructions, especially in images with repetitive structures and high-resolution details. While not leading on all benchmarks, our models offer a compelling trade-off between performance and model size, which is advantageous in practical applications where computational resources are limited and efficiency is crucial.
Future works could explore scaling our model and optimizing it for low-resolution images to improve performance on datasets like Set5 and Set14. By enhancing network depth, incorporating more advanced feature extraction techniques, and refining the balance between Transformer and Mamba components. Additionally, extending our approach to other low-level vision tasks could further demonstrate the versatility and effectiveness of hybrid architectures.
In summary, Contrast highlights the potential of hybrid architectures in advancing super-resolution performance where capturing extensive global context is essential. Our work opens new avenues for developing efficient and powerful super-resolution models that balance computational efficiency with high-quality image reconstruction, contributing to the ongoing progress in the field of low-level computer vision.